# BOPIM: Bayesian Optimization for influence maximization on temporal networks

### v2

@@ 2.3 Acquisition function
With our GP model fully defined, we can move on to the final step in the BO algorithm: the acquisition function. This function determines which seed set to evaluate our objective function at next. We want to probe areas of the search space with a high likelihood of being near the global maximum (exploitation), while also considering seed sets which have not been explored yet to prevent getting stuck in a local optimum (exploration).
For the acquisition function, we use the Expected Improvement function (Frazier,, 2018)as it is one of the most popular acquisition functions and does not require any hyper-parameter tuning. The idea is to evaluate the objective function at the value which leads to the largest expected increase compared to the current maximum. Specifically, assume that we are in $b$th iteration of the BO loop. Let $\bm{x}\in\{0,1\}^{n}$ and $\mu(\bm{x})$ and $\sigma^{2}(\bm{x})$ be the conditional mean and variance, respectively, of our GP model given the current data. By properties of the multivariate normal distribution, we have
$\mu(\bm{x})=\bm{x}^{T}\bm{\beta}+\kappa^{*}(\bm{x},{\bf X})^{T}(\gamma{\bf K}+{\bf I}_{N_{0}+b})^{-1}(\bm{y}-{\bf X}\bm{\beta})$
@@
$\sigma^{2}(\bm{x})=\sigma^{2}\{(\gamma+1)-\kappa^{*}(\bm{x},{\bf X})^{T}(\gamma{\bf K}+{\bf I}_{N_{0}+b})^{-1}\kappa^{*}(\bm{x},{\bf X})\}$
@@
In practice, we replace $\bm{\beta}$ and $\sigma^{2}$ with their corresponding posterior medians. Now, let $\Delta(\bm{x})=\mu(\bm{x})-f^{*}$ where $f^{*}=\max_{i\in\{1,\dots,N_{0}+b}\{f(\bm{x}_{i})\}$, i.e., the maximum influence spread of the seed sets we have already evaluated. Then
$EI(\bm{x})=\max\{\Delta(\bm{x}),0\}+\sigma(\bm{x})\phi\left(\frac{\Delta(\bm{x})}{\sigma(\bm{x})}\right)-|\Delta(\bm{x})|\Phi\left(\frac{\Delta(\bm{x})}{\sigma(\bm{x})}\right)$ (6)
@@
where $\phi(\cdot)$ and $\Phi(\cdot)$ are the probability density function and distribution function, respectively, of the standard normal (Frazier,, 2018). We choose the next seed set to evaluate, $\tilde{\bm{x}}$ as the seed set with largest expected improvement, i.e.,
$\tilde{\bm{x}}={\arg\max}_{\bm{x}\in\{0,1\}^{n}}EI(\bm{x}).$
@@
The astute reader will notice that we have simply traded one combinatorial optimization problem (maximizing $f(\cdot)$) for another (maximizing $EI(\cdot)$). $EI(\cdot)$, however, is much cheaper to evaluate than $f(\cdot)$ so finding its optimum will be significantly faster. Indeed, we propose a simple greedy algorithm to find the maximum of $EI(\cdot)$ with details in the Supplemental Materials. The main idea is that nodes in the current seed set are swapped with other nodes. If the switch improves the solution (larger $EI(\cdot)$), then its kept, otherwise the previous seed set is used. In this way, the algorithm makes the locally optimal decision. The algorithm continues until all nodes have been looped through and no changes are made.
@@
After solving for $\tilde{\bm{x}}=\arg\max EI(\bm{x})$, we append it and $f(\tilde{\bm{x}})$ to ${\bf X}$ and $\bm{y}$, respectively, update the covariance matrix ${\bf K}$, and re-fit the GP regression model to update the posterior distribution of $\bm{\beta}$ and $\sigma^{2}$. This step is repeated $B$ times where $B$ is the number of iterations in the BO loop. Then the optimal seed set is $\bm{x}^{*}=\arg\max_{\bm{x}\in\{\bm{x}_{1},\dots,\bm{x}_{N_{0}+B}\}}f(\bm{x})$ where this maximum is taken over the $N_{0}+B$ sampled points.
@@
Input: Objective function $f(\bm{x})$; temporal network $\mathcal{G}$; seed set size $k$; budget $B$; number of initial samples $N_{0}$;
@@
Construct ${\bf K}$ using (5)  Compute posterior of $\bm{\beta}$ and $\sigma^{2}$for  $B$ times do
       Find $\tilde{\bm{x}}=\arg\max_{\bm{x},\sum_{j}x_{j}=k}EI(\bm{x})$   Compute $y=f(\tilde{\bm{x}})$  Append observation and re-compute posterior of $\bm{\beta}$ and $\sigma^{2}$  Update ${\bf K}$
@@ 2.4 Discussion
Finding the seed set for IM on temporal networks is a cardinality-constrained, combinatorial optimization problem with non-Euclidean inputs. The challenges associated with the constraint are mitigated with our simple surrogate function (3) that is linear in the marginal contribution of each node to the influence spread. This mean function, however, implicitly assumes that there is no interaction between nodes in the seed set. From previous studies on IM, we know that this is not the case as there can be large overlap in the influence of a particular node. Instead of accounting for this property in the mean function, we explicitly incorporate the interaction between nodes via the kernel function. Our kernel function quantifies the similarity in potential seeds sets by the number of common neighbors at time $t=1$.
$\mathsf{BOPIM}$ is much faster than a standard greedy approach since the costly objective function $f(\cdot)$ only needs to be evaluated $N_{0}+B$ times, as opposed to $O(nk)$ times for a greedy algorithm. Thus, the number of evaluations of the objective function for $\mathsf{BOPIM}$ is independent of $k$. Moreover, it is trivial to adapt the algorithm to other diffusion mechanisms. In practice, the diffusion model should be chosen with great care based on the application at hand. But since the diffusion model only appears in the evaluation of the objective function, and the objective function is treated as a “black-box”, changing the diffusion model has no impact on the implementation of the method. This differs from, e.g., Dynamic Degree (Murata and Koga,, 2018), which only works for the SI model, or Erkol et al., (2020), which is based on the SIR model.
In addition to its efficiency, $\mathsf{BOPIM}$ allows for uncertainty quantification (UQ) of the total influence spread, $\sigma(S)$, and seed set, $S$. By evaluating a given seed set using the posterior distribution of $\bm{\beta}$, we obtain a complete posterior predictive distribution for the total influence spread. This is in contrast with node-ranking heuristics, which do not yield an estimated influence spread, and probabilistic approximations, which yield a point estimate but not a measure of uncertainty of the total spread. Moreover, if a researcher wants to estimate the expected influence spread for another seed set $S^{\prime}\neq S$, then the entire calculation needs to be re-run for previous methods. After $\mathsf{BOPIM}$ has been fit once, the influence spread for any seed set can be estimated immediately and the uncertainty in the optimal seed set can also easily be quantified. Please see Section 3.4 for a full discussion of UQ in the seed sets.
@@ 2.5 Hyper-parameter selection
Finally, there are several hyper-parameters that need to be chosen for $\mathsf{BOPIM}$. Choosing $\gamma$, the kernel hyper-parameter, is perhaps the most tricky. We note that $\gamma$ corresponds to the contribution of ${\bf K}$ to the overall variance of the GP regression model. If $\gamma$ is small, then most of the variance is simply coming from the random noise term, $\epsilon_{i}$, but if $\gamma$ is large, then the variance in the objective function evaluations is largely driven by our kernel function. There are many different ways that we could choose $\gamma$. If we want a fully Bayesian approach, then we could endow $\gamma$ with a prior and update its posterior distribution during the MCMC routine. In practice, however, we found that this did not lead to very good performance and required adding a Metropolis-Hastings step to an otherwise entirely Gibbs sampler. Another choice is to fix $\gamma$ at its maximum likelihood estimate (MLE). In our experiments, this choice proved to perform better than the fully Bayesian approach, but the MLE estimate of $\gamma$ was consistently close to 1. Indeed, if instead of finding the MLE for $\gamma$ after each new observation we simply set $\gamma=1$, then this yielded the best performance. This choice corresponds to equal contribution from ${\bf K}$ and the random error term.
There are several other parameters that must be chosen in our algorithm. For the number of initial samples $N_{0}$ and BO loop iterations $B$, if these values are large, then the surrogate function will fit the true function well, but at a high computational cost. We found that setting $N_{0}=5$ and $B=25$ led to a good balance of fit and speed. To show that the BO iterations improve the solution quality, we report a progress curve in Figure 1 for a single run of the $\mathsf{BOPIM}$ algorithm. Please see Section 3 for a description of the network. We can clearly see that as the number of BO iterations increases, the total influence spread also increases. Indeed, for this particular example, the solution quality (percentage of influenced nodes) increases by 18 % during the BO loop, which corresponds to a relative increase of almost 100%.
@@
For computing the posterior of $\bm{\beta}$ and $\sigma^{2}$, the number of MCMC samples also must be selected. Again, more MCMC samples lead to better estimates of the posterior distribution but also longer run-times. Since we only use the median of the posterior distribution, we can keep the number of MCMC samples relatively small to increase speed. We recommend 2,000 iterations with the first 500 used for burn-in. For hyper-parameters in the Horseshoe prior, we simply use the default settings from the original papers. Lastly, for the number of MC simulations for the objective function calculation, it is important that this output is accurate so we recommend 1,000 iterations. Unless otherwise noted, we use these settings for all experiments in Section 3.
@@ 3 Experiments
We now study the performance of the proposed method with numerical experiments. All experiments were performed in Python and the code for Algorithm 1 is available on GitHub: https://github.com/eyanchenko/BOPIM.
@@ 3.1 Data
The following experiments are conducted on four real-world networks. Reality ($n=64$, $m=722$) (Eagle and Pentland,, 2006), Hospital ($n=75$, $m=1139$) (Vanhems et al.,, 2013), Bluetooth ($n=136$, $m=5949$) (Kotz and Henderson,, 2004) and Conference 2 ($n=199$, $m=1550$) (Chaintreau et al.,, 2007) are all proximity networks where an edge exists if two people are close to each other. In order to discretize the continuous appearing and disappearing of edges, each network is aggregated into $T$ temporal snapshots of equal duration. We note that these networks are relatively small due to the slow speed of the greedy algorithm. Therefore, in the Supplemental Materials, we also conduct an experiment using the proposed method on the DNC email network (Rossi and Ahmed,, 2015) with $n=1891$ nodes to demonstrate its scalability.
3.2 Surrogate function validation
3.2.1 Settings
In the first set of simulations, we quantify the fit of the surrogate model to the true objective function $f(\cdot)$. Towards this end, we first run Algorithm 1 with $N_{0}=5$ and $B=25$ and save the final posterior distribution of $\bm{\beta}$. Using this posterior distribution, we then compute out-of-sample influence spread predictions on $N=100$ seed sets and compare with the values of the true objective function. If the surrogate model fits well, then its predicted influence spread will be close to the true influence spread. We quantify this similarity using the mean absolute prediction error (MAPE). If $\widehat{\bm{\beta}}$ is the posterior median of $\bm{\beta}$, and $\bm{x}^{*}$ corresponds to the new seed set, then
$\hat{y}=(\bm{x}^{*})^{T}\hat{\bm{\beta}}+\kappa(\bm{x}^{*},{\bf X})^{T}(\gamma{\bf K}+{\bf I}_{N_{0}+b})^{-1}(\bm{y}-{\bf X}\hat{\bm{\beta}})$
where $\kappa(\bm{x}^{*},{\bf X})_{j}=\kappa(\bm{x}^{*},\bm{x}_{j})$ for $j=1,\dots,N_{0}+B$. Recall that $\kappa(\bm{x}^{*},\bm{x}_{j})$ is the Jaccard coefficient between the neighbors of the seed set corresponding to $\bm{x}_{j}$ and $\bm{x}^{*}$. The MAPE is computed as
$\frac{1}{N}\sum_{i=1}^{N}|\hat{y}_{i}-y_{i}|$
We compute the MAPE for each of the four data sets. We fix $\lambda=0.05,0.05,0.01,0.01$, and $k=3,4,6,10$ for the Reality, Hospital, Bluetooth and Conference 2 datasets, respectively. $k$ was chosen as approximately 5% of the total number of nodes for each network. For the out-of-sample predictions, we consider two node sampling schemes. One samples nodes at random while the other samples according to their degree on the aggregate network, as in $\mathsf{BOPIM}$. We expect the model to have a closer fit when the test points are sampled in the same way as the training points. Results are averaged over 25 Monte Carlo (MC) replications.
3.2.2 Results
The results are in Table 1. Note that the MAPE results are interpreted as follows: for the Conference 2 network with degree-based sampling of seed sets, for example, the MAPE value of 1.23 means that, on average, the estimated influence spread of the surrogate function (3) is within about one node of the true influence spread. We can see that for all networks and both sampling schemes, the MAPE values are quite small. Additionally, the degree-based sampling always leads to smaller MAPE values, which is to be expected since this is how the $\mathsf{BOPIM}$ algorithm samples points for the initial fitting stage. Indeed, the surrogate model fits the true objective function more closely when the seed nodes have a large degree, which is where we hypothesize that the global optimum lies. These results give use confidence that our surrogate function is doing an adequate job of modeling the true objective function.
Dataset Sampling MAPE (se)
Reality Random 3.60 (0.19)
Degree 3.22 (0.11)
Hospital Random 4.88 (0.16)
Degree 3.49 (0.12)
Bluetooth Random 2.78 (0.08)
Degree 1.77 (0.04)
Conference 2 Random 7.45 (0.11)
Degree 1.23 (0.05)
3.3 Influence maximization
For the second set of experiments, we compare the method’s performance on the IM task.
3.3.1 Settings

### v4

@@ 2.4 Acquisition function
With our GP model fully defined, we can move on to the final step in the BO algorithm: the acquisition function. This function determines which seed set to evaluate the objective function at next. We want to probe areas of the search space with a high likelihood of being near the global maximum (exploitation), while also considering seed sets which have not been explored yet to prevent getting stuck in a local optimum (exploration). Moreover, this function will also need to account for the fact that we observe the influence spreads with noise.
We prefer the Expected Improvement acquisition function (Močkus,, 1975; Jones et al.,, 1998; Frazier,, 2018) as it is one of the most popular acquisition functions and does not require any hyper-parameter tuning. The idea is to evaluate the objective function at the value which leads to the largest expected increase compared to the current maximum. Specifically, assume that we are in $b$-th iteration of the BO loop. Let $\bm{x}\in\{0,1\}^{n}$ and $\tilde{\mu}^{(b)}(\bm{x})$ and ${\tilde{\sigma}}^{(b)}(\bm{x})$ be the conditional mean and standard deviation, respectively, of our GP model given the current data, where we explicitly emphasize the dependence on the iteration with the super-script in the notation. By properties of the multivariate normal distribution, we have
$\tilde{\mu}^{(b)}(\bm{x})=\beta_{0}^{(b)}{\bf 1}_{N_{0}+b}+\kappa^{*}(\bm{x},{\bf X})^{T}(\gamma{\bf K}+{\bf I}_{N_{0}+b})^{-1}(\bm{y}-\beta_{0}^{(b)}{\bf 1}_{N_{0}+b})$
@@
${\tilde{\sigma}}^{(b)}(\bm{x})={\sigma}^{(b)}\{(\gamma+1)-\kappa^{*}(\bm{x},{\bf X})^{T}(\gamma{\bf K}+{\bf I}_{N_{0}+b})^{-1}\kappa^{*}(\bm{x},{\bf X})\}^{1/2}$
@@
In practice, we replace $\beta_{0}^{(b)}$ and ${\sigma}^{(b)}$ with their corresponding posterior medians. If our observations were noiseless, we would let $\Delta^{(b)}(\bm{x})=\tilde{\mu}^{(b)}(\bm{x})-{f^{(b)}}^{*}$ where ${f^{(b)}}^{*}=\max_{i\in\{1,\dots,N_{0}+b}\{y_{i}\}$, i.e., the maximum influence spread of the seed sets we have already evaluated. Then
$EI^{(b)}(\bm{x})=\max\{\Delta^{(b)}(\bm{x}),0\}+\tilde{\sigma}^{(b)}(\bm{x})\phi\left(\frac{\Delta^{(b)}(\bm{x})}{\tilde{\sigma}^{(b)}(\bm{x})}\right)-|\Delta^{(b)}(\bm{x})|\Phi\left(\frac{\Delta^{(b)}(\bm{x})}{\tilde{\sigma}^{(b)}(\bm{x})}\right)$ (5)
where $\phi(\cdot)$ and $\Phi(\cdot)$ are the probability density function and distribution function, respectively, of the standard normal (Frazier,, 2018). Recall, however, that the influence spread, $y_{i}$, is observed with Gaussian noise $\varepsilon_{i}$, making ${f^{(b)}}^{*}$ a (potentially) poor and non-robust estimate of the global maximum, particularly if it was observed with large noise. Following Vazquez et al., (2008) and Picheny et al., (2013) , we replace ${f^{(b)}}^{*}$ with a “plug-in” estimate of the global maximum; namely, the maximum of the conditional mean function, $\tilde{\mu}^{(b)}(\cdot)$, evaluated at the previously sampled points, i.e.,
${f^{(b)}}^{*}=\max_{i\in\{1,\dots,N_{0}+b\}}\tilde{\mu}^{(b)}(\bm{x}_{i}).$
We are still not quite finished as the acquisition function in (5) implies that we observe the next observation deterministically. To account for the noise in the ensuing observations, we included a multiplicative term as in Huang et al., 2006a ; Huang et al., 2006b , which penalizes sampling a new seed set where the conditional standard deviation, ${\tilde{\sigma}}^{(b)}(\cdot)$ is small. This is a sensible choice, as a small variance implies we already know the influence spread of this seed set with fairly high certainty, encouraging exploring elsewhere in the input space. The final (augmented) expected improvement (AEI) acquisition function becomes
$AEI^{(b)}(\bm{x})=EI^{(b)}(\bm{x})\times\left(1-\frac{{\sigma^{2}}^{(b)}}{\sqrt{\{\tilde{\sigma}^{(b)}(\bm{x})\}^{2}+{\sigma^{2}}^{(b)}}}\right),$ (6)
@@
and we choose the next seed set to evaluate, $\tilde{\bm{x}}$ as the seed set with largest expected improvement, i.e.,
$\tilde{\bm{x}}={\arg\max}_{\bm{x}\in\{0,1\}^{n}}AEI^{(b)}(\bm{x}).$
@@
The astute reader will notice that we have simply traded one combinatorial optimization problem (maximizing $f(\cdot)$ for maximizing $AEI(\cdot)$). $AEI(\cdot)$, however, is much cheaper to evaluate than $f(\cdot)$ so finding its optimum will be significantly faster. Indeed, we propose a simple greedy algorithm to find the maximum of $AEI(\cdot)$ with details in the Supplemental Materials. The main idea of this algorithm is that nodes in the current seed set are swapped with other nodes. If the switch improves the solution (larger $AEI(\cdot)$), then its kept, otherwise the previous seed set is used. In this way, the algorithm makes the locally optimal decision. The algorithm continues until all nodes have been looped through and no changes are made. Please see the Supplemental Materials for a connection between the proposed greedy algorithm and well-known Trust region (TR) framework (e.g., Eriksson et al.,, 2019; Wan et al.,, 2021; Papenmeier et al.,, 2023)
@@
After solving for $\tilde{\bm{x}}=\arg\max AEI^{(b)}(\bm{x})$, we append it and its evaluated influence spread to ${\bf X}$ and $\bm{y}$, respectively, update the covariance matrix ${\bf K}$, and re-fit the GP regression model to update the posterior distribution of $\beta_{0}$ and $\sigma^{2}$. This step is repeated $B$ times where $B$ is the number of iterations in the BO loop. Then the optimal seed set is
$\bm{x}^{*}=\arg\max_{\bm{x}\in\{\bm{x}_{1},\dots,\bm{x}_{N_{0}+B}\}}\tilde{\mu}^{(B)}(\bm{x}),$
where this maximum is taken over the $N_{0}+B$ sampled points and depends on the final posterior distribution of $\beta_{0}$ and $\sigma^{2}$.
@@
Input: Objective function $f(\cdot)$; temporal network $\mathcal{G}$; seed set size $k$; budget $B$; number of initial samples $N_{0}$;
@@
Construct ${\bf K}$ using (3) or (4)   Compute posterior of $\beta_{0}$ and $\sigma^{2}$for  $B$ times do
       Find $\tilde{\bm{x}}=\arg\max_{\bm{x},\sum_{j}x_{j}=k}AEI(\bm{x})$   Compute $y=f(\tilde{\bm{x}})$  Append observation and re-compute posterior of $\beta_{0}$ and $\sigma^{2}$  Update ${\bf K}$
@@ 2.5 Discussion
Finding the seed set for IM on temporal networks is a cardinality-constrained, combinatorial optimization problem with non-Euclidean inputs. We primarily address these challenges with the kernel function and acquisition function. Both of these components of our BO algorithm are specifically tailored to handle the unique challenges posed by IM.
$\mathsf{BOPIM}$ is much faster than a standard greedy approach since we only need to obtain $N_{0}+B$ realizations of the costly objective function, as opposed to $O(nk)$ times for a greedy algorithm. Thus, the number of evaluations of the objective function for $\mathsf{BOPIM}$ is independent of $k$. Moreover, it is trivial to adapt the algorithm to other diffusion mechanisms. In practice, the diffusion model should be chosen with great care based on the application at hand. But since the diffusion model only appears in the evaluation of the objective function, and the objective function is treated as a “black-box”, changing the diffusion model has no impact on the implementation of the method. This differs from, e.g., Dynamic Degree (Murata and Koga,, 2018), which only works for the SI model, or Erkol et al., (2020), which is based on the SIR model.
One of the main advantages of the BO framework is that it easily allows for quantifying uncertainty. We discuss this in more detail in Section 4, but briefly anticipate that discussion. By evaluating a given seed set using the posterior distribution of $\beta_{0}$ and $\sigma^{2}$, we obtain a complete posterior predictive distribution for the total influence spread. This is in contrast with node-ranking heuristics, which do not yield an estimated influence spread, and probabilistic approximations, which yield a point estimate but not a measure of uncertainty of the total spread. Moreover, if a researcher wants to estimate the expected influence spread for another seed set $S^{\prime}\neq S$, then the entire calculation needs to be re-run for previous methods. After $\mathsf{BOPIM}$ has been fit once, the influence spread for any seed set can be estimated immediately and the uncertainty in the optimal seed set can also easily be quantified.
@@ 2.6 Hyper-parameter selection
Finally, there are several hyper-parameters that need to be chosen for $\mathsf{BOPIM}$. Choosing $\gamma$, the kernel hyper-parameter, is perhaps the most tricky. We note that $\gamma$ corresponds to the contribution of ${\bf K}$ to the overall variance of the GP regression model. If $\gamma$ is small, then most of the variance is simply coming from the random noise term, $\epsilon_{i}$, but if $\gamma$ is large, then the variance in the objective function evaluations is largely driven by our kernel function. There are many different ways that we could choose $\gamma$. If we want a fully Bayesian approach, then we could endow $\gamma$ with a prior and update its posterior distribution during the MCMC routine. In practice, however, we found that this did not lead to good performance and required adding a Metropolis-Hastings step to an otherwise entirely Gibbs sampler. Another choice is to fix $\gamma$ at its maximum likelihood estimate (MLE). In our experiments, this choice proved to perform better than the fully Bayesian approach, but the MLE estimate of $\gamma$ was consistently close to 1. Indeed, instead of finding the MLE for $\gamma$ after each new observation, simply setting $\gamma=1$ yielded the best performance. This choice corresponds to equal contribution from ${\bf K}$ and the random error term.
There are several other parameters that must be chosen in our algorithm. For the number of initial samples $N_{0}$ and BO loop iterations $B$, if these values are large, then the surrogate function will fit the true function well, but at a high computational cost. We found that setting $N_{0}=5$ and $B=20$ led to a good balance of fit and speed. To show that the BO iterations improve the solution quality, we report a progress curve in Figure 1 for a single run of the $\mathsf{BOPIM}$ algorithm. Please see Section 3 for a description of the network. As the number of BO iterations increases, the total influence spread also increases. Indeed, for this particular example, the solution quality (proportion of nodes influenced) increases by over 16% during the BO loop, which corresponds to a relative increase of almost 70%.
@@
We use a standard Gibbs sampler to estimate the posterior distribution of $\beta_{0}$ and $\sigma^{2}$, meaning the number of MCMC samples must also be chosen. More MCMC samples leads to better estimates of the parameters, but also longer run-times. Since we only use the median of the posterior distribution, we can keep the number of MCMC samples relatively small to increase speed. We recommend 2500 iterations with the first 500 used for burn-in. Lastly, for the number of MC simulations for the objective function realization, it is important that this output is accurate so we recommend 1,000 iterations. Unless otherwise noted, we use these settings for all experiments in Section 3.
@@ 3 Experiments
We now study the performance of the proposed method with numerical experiments. All timed experiments were performed in Python on a 2024 M4 Mac Mini with 16 GB of memory and 9 cores in use. The code for Algorithm 1 is available on GitHub: https://github.com/eyanchenko/BOPIM.
@@ 3.1 Data
The following experiments are conducted on four real-world networks. Reality ($n=64$, $m=722$) (Eagle and Pentland,, 2006), Hospital ($n=75$, $m=1139$) (Vanhems et al.,, 2013), Bluetooth ($n=136$, $m=5949$) (Kotz and Henderson,, 2004) and Conference 2 ($n=199$, $m=1550$) (Chaintreau et al.,, 2007) are all proximity networks where an edge exists if two people are close to each other. In order to discretize the continuous appearing and disappearing of edges, each network is aggregated into $T$ temporal snapshots of equal duration. We note that these networks are relatively small due to the slow speed of the greedy algorithm. Therefore, in the Supplemental Materials, we also conduct an experiment using the proposed method on the DNC email network (Rossi and Ahmed,, 2015) with $n=1891$ nodes to demonstrate its scalability. Additionally in the Supplemental Materials, we also quantify the fit of the surrogate model to the true objective function.
3.2 Settings